SYLLABUS FOR MATH 141, FALL 2010

Monday, Wednesday, Friday 11:00 a.m. - 11:50 a.m. Classroom: ARM 0131

Prerequisites: A grade of C or better in MATH 140 or equivalent

Text: Calculus (Sixth Edition) by R. Ellis and D. Gulick, ISBN 978-0-7593-1379-2 or 0759315299 (Required)

Instructor: Wojciech Czaja Teaching Assistants:

Mr. Bryant Angelos email: bangelos (at) math (dot) umd (dot) edu

sections 251 and 261

Mr. Michael Kreisel email: mkreisel (at) math (dot) umd (dot) edu

sections 212 and 222

Mr. Timothy Mercure email: mercure (at) math (dot) umd (dot) edu

sections 231 and 241

Mr. Nathaniel Monson email: nmonson1 (at) math (dot) umd (dot) edu sections 211 and 221

Office: MTH 4406 Office Hours: Monday, Wednesday, Friday 10:00 a.m. - 10:50 a.m. or by appointment Phone: (301) 405 - 5106 Email: wojtek (at) math (dot) umd (dot) edu (Monday - Friday)

My Web Page: www.math.umd.edu/∼wojtek

Class Web Page: www.math-users.umd.edu/∼wojtek/141 10.htm

Exams: There will be 3 midterm exams during the semester, each for the total of 100 points. They are scheduled on September 22, October 22, and November 22. The lowest score (including absences) is dropped.

The final exam is scheduled for December 13, 1:30-3:30 PM (rooms vary by TA). It is worth 200 points.

Calculators: You will need a programmable graphics calculator for some homework and section quizzes. You will need to put two programs on the calculator. You are responsible for knowing how to operate your calculator, but for some Texas Instruments graphics calculators (TI-82,83, ...) the two programs are on the course page.

Calculators and any other electronic devices (including cell phones) are NOT allowed on midterm tests, nor on the final.

Quizzes: 7 quizzes will be given during the semester. Each is worth 20 points. The lowest 2 scores (including absences) will be dropped. Thus, the total weight of quizzes is 100 points. The quizzes will be given during the discussion sections on the following days: 09/07, 10/05, 10/19, 11/04, 11/18, 12/02, 12/09.

Homework: There will be 27 Webassign (online) homeworks during the semester. Each homework is worth 5 points. The lowest scores will be dropped, so that the total weight of homeworks is 100 points.

Grading: The maximum point total is 600 points. The total used to calculate the final grade is the minimum of the two numbers: 600 points and the highest score in the class. The setting of letter grades will be based on the number of points and will be not worse than: 50% - D, 60% - C, 75% - B, 90% - A.

Academic integrity: The University makes me remind you about its academic integrity policies (http://www.shc.umd.edu/code.html). So I do. Nobody, however, has to remind me that part of my job is to make sure that these policies are obeyed.

Attendance and absences: You are responsible for the material covered in class, whether you attend or not. You are also responsible for the announcements made during class; they may include changes in the syllabus. The instructor will adhere strictly to the official university policy on makeup exams. In particular, it is the policy of the University to excuse the absences of students that result from the following causes: illness of the student, or illness of a dependent; religious observance (where the nature of the observance prevents the student from being present during the class period); participation in university activities at the request of University authorities; and compelling circumstance beyond the student’s control. Students claiming excused absence must apply in writing and furnish documentary support for their assertion that absence resulted from one of these causes. Moreover, foreseeable absences (such as those resulting from religious holidays or participation in university-sponsored events) must be submitted in writing to the instructor by Sept.

15. (See also http://www.testudo.umd.edu/soc/atedasse.html)

Disabilities: If you have a disability disallowing you to test under the usual time or in-class conditions, you may contact the office of Disabled Students Services (DSS) in Shoemaker. (Please let me know by Sept. 15 if you think you may require these services.) If they assess you as meriting private conditions and/or extra time, then you may arrange to take your tests at DSS, with extra time as they indicate. You must arrange this well in advance of a test (in particular: no retakes). Remember: I need to sign the document.

Emergency closures: In case of an emergency that closes the University for an extended period of time (for example, due to inclement weather), be sure to access your email for instructions from me. Also check the University’s home page or call 301-405-SNOW for snow closure information. Should any classes or exams be cancelled, please check the class schedule page for updated schedule information.

Course evaluations: Your participation in the evaluation of courses through CourseEvalUM is important to us, and helps improve teaching and learning at the University. CourseEvalUM will be open for you to complete your evaluations for the Fall semester courses between Tuesday, November 30, and Sunday, December 12. Please provide feedback on the course, the professor, and the TA by Sunday, December 12, 2010. Evaluations are anonymous and will not be available to faculty and TAs until next semester, so they cannot possibly affect your grade.

You can go directly to the website (www.courseevalum.umd.edu) to complete your evaluations starting November 30. By completing all of your evaluations each semester, you will have the privilege of accessing the summary reports for thousands of courses online at Testudo.

When necessary, the syllabus will be adjusted and the changes will be announced in class.

OUTLINE OF MATERIAL

08/30 Introduction, 6.1 09/01 6.1, 6.2 09/03 6.3 09/06 Labor day 09/08 6.4 09/10 6.5 09/13 6.7 09/15 6.8 09/17 7.1 09/20 Review for Exam 1 09/22 Exam 1 09/24 7.2, 7.3 09/27 7.5 09/29 7.6 10/01 7.7 10/04 7.8 10/06 8.1 10/08 8.2 10/11 8.3 10/13 8.4 10/15 8.6 10/18 8.7 10/20 Review for Exam 2 10/22 Exam 2 10/25 9.1 10/27 9.2 10/29 9.3 11/01 9.4 11/03 9.4 11/05 9.5 11/08 9.5, 9.6 11/10 9.7 11/12 9.8 11/15 9.9 11/17 9.9, 9.10 11/19 Review for Exam 3 11/22 Exam 3 11/24 Discussion 11/26 Thanksgiving 11/29 Complex numbers 12/01 Complex series 12/03 Complex plane 12/06 10.1 12/08 10.2 12/10 Review for the final exam